DFR Telecoms Diary
1 January :
9 January : Sam, Martin, Ray, Charles, Peter and I were in. We spent the day starting on the big clean up of the caboosh. We have quite a pile of gear that will go to the next THG swapmeet in April. Peter and Ray checked out a Panasonic PABX which we think is going to the railway at Blaenavon.
16 January : Discussions were held about moving line plant recording from John to Martin, fault logging and lineside visit recording. I spent my day jumpering up three of the new relay sets, the transformer bridge circuit for Parkend North, followed by the two low back emf repeaters in the asterisk incoming circuits. They all worked though I did have a minor fault on one relay set where a resistor was touching the earthed frame. . Ray went over the clock circuitry with Sam. Peter and Martin had a look at the cabling in the main building Severn and Wye room. Sam and Peter carried on with altering the bench in our office in preparation for the move of the data cabinet to the corner.
19 January : I went in to reinstall the two first selectors. One had a fault which turned out to be a dis wire caused by me tugging on the U points when I pulled them out of the frame. They are both now working. I took number 4 home to demodify. I expect this one to be much more difficult as it has been more extensively modified. I came back to see Sam at Norchard. He and Roger want me to move an auto auto with regenerator relay set from Parkend to Norchard but I am pretty sure that it does not exist. It was discussed but never built. I have the relay can and regenerator at home available for the build. It's probably a good idea for both Norchard to Asterisk junctions to have regenerators as this will clean up any dial pulses from faulty dials. No chance of doing this at Parkend also, but I guess that all calls to the digital network can be connected at Norchard only. Calls from the Parkend digital network to the Strowger can be handled at Parkend without any problem as the low back EMF relay set has been built and just needs cabling in.
19 January : It is proposed to move the line plant record keeping from John Bathgate to Martin Lunn. The following points outline the procedures we intend to adopt.
20 January :
23 January : It's still a clearing up time at the railway. Rick and Martin were in the exchange sorting junk for the sale table or the skip. Charles had a look at the faulty light in the exchange but it's a LED tube in a modified fluorescent fitting so he is still thinking about it. Sam and Peter went surveying the situation at the Junction and at Parkend may have come across a BT jointer who could help with the external joint required at DP1E. I went to Parkend to reinstall the first selector I had demodified. Then I attempted to wire the two wire repeater relay sets to the IDF but was diverted by a fault in Parkend North exchange, then a fault in the U points of the relay set and faults in the cradles holding the associated incoming selectors. By the time I had sorted out what I needed to do it was time to go home.
30 January : I did nothing as the existing NCD to ASK junctions are not working. I did not want to add new untested gear to faulty junctions. Rick is tidying and sorting spares in Norchard exchange. Charles continues with his light saga. Sam tried to get the Asterisk back to health but failed. Not a very productive day. But it was freezing outside! Sam has bought a couple of Rasberry Pies using our fund in order to initially set up a demo digital system.
February 2019 :
6 February : Peter, Martin and Charles looked at records, Alan and Sam experimented with a demo asterisk system. I went to Parkend and cabled, terminated and jumpered the two wire repeaters into circuit.
13 February : Sam and Alan repaired an S&T circuit at Parkend, Rick and Martin continued to work in Norchard exchange while I went to Parkend to make a start on getting rid of the hard to reach IDF blocks.
20 February : Some records were checked first. Rick and Martin are continuing with getting Norchard overhauled. This will take a very long time as they are really going to town on getting the equipment, cleaned and properly adjusted. Sam and Ray are having a go at getting the new PA speakers connected up. I have now got all five first selectors at Parkend demodified and working as standard. I checked the junctions from Parkend to find Junction 1 to Norchard dis out and very noisy. It's been busied but needs attention next time.
27 February : Rick and Martin found the dis in Parkend Junction 3 at DP3Y. Bad connection again. They went back to maintenance at Norchard. Charles had a go at sorting the cable for sale to Bendall's and getting connectors etc cut off. I went to Parkend and got the three circuit auto auto relay set off the old IDF block and onto the new block. Much easier to jumper now.
February 2019 :
6 March : Rick and Martin are continuing to do the maintenance in Norchard. Sam and Peter checked the Lydney Junction SPTs and had to correct some line plant errors to get everything working. The PA is not working at the Junction, also probably due to line plant rearrangement errors. Ray and I worked on a fault at Parkend but found nothing conclusive, every now and then a selector drops out, but nothing is consistent. I have now cleared one of the old IDF blocks, the other has only a few miscellaneous services on it and another day should see the old IDF cleared and recovered. Then I will have to tidy all the wiring forms on the rack as they have all been cut open for wiring rearrangements.
9 March : Roger and Sam met up at Parkend, before decanting to Lydney Junction to investigate the broken PA in the station building. When we looked into the DP inside (1B), we discovered that it was significantly different than the records! We wired it to how it should be, toned out the circuits from the MDF, and soon had the Beatles pounding out across Lydney from the speakers!
13 March : Our day started with a discussion on the regular maintenance of "safety critical" telecoms gear. Even with S&T involved this did not lead to any precise definition. Sam is to write an instruction on such maintenance. I am to put together a training package on our concentrators as I have ended up being the only person who is competent to maintain them. Sam and Alan then went on to complete the annual tests of all Norchard phones and to make some necessary repairs. Martin, Rick and Charles returned to Norchard exchange maintenance. I returned to Parkend where I shifted the incoming final selectors onto the new IDF block.
20 March : The first Wednesday of the running season but we could only muster the 08 for the train. The point at Middle Forge has been damaged and has prevented the steam locos from moving out onto the line. I finished moving the circuits off the old IDF blocks at Parkend and got the blocks recovered. That makes the rack look tidier and easier to jumper on. I still have to wire up the new direct telephone relay set and connect the common services it needs. Then everything needs fully checking and when all is working all the cable forms need relacing. Sam and Charles went looking for a cable route between DPs 3a and 3E. This involved in getting under the cafe building. Martin, Ray and Alan were sorting spare stores between scrapping, keeping and selling. Peter is starting to look at how best to relieve and reorganise DP3E and Norchard box. We also have a request to provide a counter for the number of people crossing the footbridge at St Mary's, Sam will work on the bridge end and I will provide the counter in the exchange at Norchard. We have a further request to provide phones in the two new cabins at the rear of the workshop, one is a canteen and the other will be a record store.
27 March : Rick and Martin chased the faulty Parkend junction yet again. They think the fault may lie within a joint. Charles and I added an extra meter circuit to our junction call counters and jumpered it back to the redundant data pair to Lydney Junction. It is now ready for Sam to fit a counter sender at the St Mary's footbridge. Peter was in for a short time collecting information about the cables and circuits around the Norchard signal box.
April 2019 :
3 April : Sam, Peter and Charles investigated the work necessary to provide the footfall counter at the bridge. The cabinet for DP3M needs to be replaced by one of the "big" railway's telecoms cabinet in order for all the pairs and the proposed twenty pair to the signal box to be terminated. Unfortunately Sam cannot proceed too far with his counter as currently there are no mains to the bridge. Rick and Martin went back on the Norchard to Whitebrook cables trying to find the intermittent "dis" on PKD to NCD outgoing 1. By lunchtime the circuit was working again but they were not sure why. I cleared a fault on NCD to PKD outgoing 1 (no transmission, even though dialling was fine). It was a worn residual on the relay set C relay letting the relay stick on most of the time. I wired out the direct call relay set at Parkend and got the common services connected. Now it's time to find out what all the spare wires are and then get the cable forms laced back up.
10 April : Sam spent the day with a representative of the Avon Valley Railway who has been charged with the provision of signalling on his railway. Rick, Martin and Peter vanished off looking at faults in the Parkend area. They spent part of the day collating DP distances from Parkend exchange. I installed the second auto auto with regenerator relay set at Norchard and then got them jumpered to level 4 and out to the MDF. A lot of rejumpering and labelling required. The regen in the existing set was faulty and needed lubricating , general readjusting and the output speed reset. These two relay sets will provide the outgoing link to the digital network. The incoming link already exists.
14 April : Sam, Peter and I attended the THG swapmeet at Avoncroft. A profitable day for our telecoms fund. Our net profit was £208.50, however we still have some expenses to pay. A good day.
15 April : Peter met Sam at Parkend exchange as some measurements were missing for DP1 and DP1A. The missing measurements will be added to a spreadsheet which will make cable length calculation much easier for planning the cable routes for the required SPT's connected to the signal box. The only measurements now required are from DP 1H south to Whitecroft. Whilst at Parkend they also worked out an entry to the exchange from the new troughing, a tee and two straight sections are all that is required and can hopefully be delivered on Wednesday.
24 April : Alan and Peter went to the Junction to work on DP 1 at the south end of the platform. A new back board was fitted in the S+T cabinet and DP 1 is now fixed in place waiting for the provision of the new 20 Pr cable to the station building. After lunch Charles and Peter went to Parkend signal box and recovered five lengths of six wire originally destined for the S+T relay rack, this will now be replaced with a 20pr + Earth cable recovered from OP3 at Norchard. The cable will then be connected to a box conn 254. The recovered six wire cables will find a new home in Norchard signal box. Sam, Charles and Martin chased a fault on an S+T circuit at Parkend which, as usual, in the end proved to be faulty S+T equipment with the cable just fine. Sam and Martin repaired the faulty S+T indicating instrument and have presented it to S+T for refitting and testing. I continued to work at Parkend. I think all the wiring has been completed and I spent the day relacing all the cable on the main rack.
30 April : Sam and Peter met at the Jasmin cafe for lunch and to discuss the the Norchard - Asterisk gateway. Sam and Peter proceeded to the railway to get the Supermicro !U short server case from the loft. Whilst Sam was in the loft he recovered DP 10, an old fashion 4 wire cable terminated on B151 & 152. Peter recovered the MDF cabling for the scrap bin. The brand new, in the box server case, proved to be obstinate in giving up 1 of the case screws. Whilst at Norchard Peter and Sam indulged in a little retail therapy for men and surfed e-bay. They purchased a mini-itx motherboard for the new gateway, at the staggering cost of £30 including memory and an embedded CPU. They decided to go to Peter's to extract the stubborn screw. When they final removed the screw using a No 1 screwdriver in a cross point screw it revealed sadly that the SuperMicro case can only be fitted with a SuperMicro motherboard, cheapest on e-bay £229. Plan B. Peter had a dig around and came up with a couple of suitable VESA type cases, these are neat and compact and should be able to be fitted in the exchange without too much trouble.
1 May : Charles and Alan laid in a 10Pr + Earth cable to the S+T relay rack and left a suitable length tail ready for David Collins to provide details of where he wishes the Box Con 254 to be fitted. Once the cable was laid in Peter terminated it on the S+T part of the MDF. Sam has had a new telecoms cabinet delivered to PKD which will be installed next to the newly provided S+T cabinets at the ground frame. Alan, Charles and Peter went walk about and measured the troughing distance from the end of the station fence to the new cabinet. Martin and Rick went to check signal post 13 - reported faulty. The dropwire had been severed through in two places probably by the tamper machine. They had to renew and reroute the dropwire from signal post 13 through to DP3G. All working fine now. After a late lunch they all had an in depth discussion about DP3E.
5 May : Peter met Sam at 12:00 for lunch in the railway cafe. What no Sunday dinners, a disappointment. After lunch Peter went to the caboosh to sort out stores for up and coming projects. This consisted of finding all the necessary box connections, krone strips and strips designations. Once assembled the box conns were placed in the relevant stores container. along with the double back boxes. Faceplate and modules where possible were added to the stores boxes. Whilst we have an abundance of LJU 2/2A and RJ45 modules we have run out of LJU 2/1A modules so some will need to sourced. The majority of stores for the junction station, new Norchard DP 3E, new Norchard crew room and Whitecroft station have been picked, this is of course with the exception of new external cable which will have to be ordered from Qing's. Peter took one of the 8064 blocks home to measure it with a view to designing a new wall mounting bracket.
8 May : Ray finally put to bed the calls dropping out when a final selector was seized at Parkend,. It was an intermittent HR on final selector 4 B relay 8 and 9 contacts which provided a holding earth back through the selectors. Cleaned and checked B relay contacts and did much testing on all selectors. All OK now. Rick worked in Norchard Exchange on maintenance - mainly wipers and cords. Sam and Charles were with John Clark looking and talking about electrical safety at Norchard. The afternoon was spent tidying and cleaning our room ready for the Thomas weekend.
15 May : Peter and Allan were surveying at Lydney Junction. Rick, Ray and Charles travelled on a round trip looking for installed equipment/damage. An incorrectly labelled phone (in the No. 1 key accessed box at Parkend) had the label removed, no fault found there.
16 May : The group went to lunch at the George at St Briavels. Sam, Rick and Janet, Charles and Alison, Ray and June, John and Jean attended a very good meal.
22 May : Rick and Martin spent their day tidying the office. We actually have a bit of floor space now. Sam and I chased a fault on the PA. We decided that if the PA played music, then when 523 in the shop was rung, the ringing could be heard in the background over the music. In the end we decided the overhearing was somewhere in DP17 or the cable to it. We shifted the music to OP2 in the office and this got rid of the overhearing. Sam also moved the music controller to the back wall next to OP2. Then he was called to do Duty Officer and vanished off on the 08 going south as it replaced a failed 5541. We still need to check that music will still play.
Current Value of our Fund
Return to the Diary Menu